# Fairness-Accuracy Trade-Offs: A Causal Perspective

### v1

@@ 1.1 Preliminaries
We use the language of structural causal models (SCMs) (Pearl, 2000). An SCM is a tuple $\mathcal{M}:=\langle V,U,\mathcal{F},P(u)\rangle$ , where $V$, $U$ are sets of endogenous (observable) and exogenous (latent) variables, respectively, $\mathcal{F}$ is a set of functions $f_{V_{i}}$, one for each $V_{i}\in V$, where $V_{i}\leftarrow f_{V_{i}}(\mathrm{pa}(V_{i}),U_{V_{i}})$ for some $\mathrm{pa}(V_{i})\subseteq V$ and $U_{V_{i}}\subseteq U$. $P(u)$ is a strictly positive probability measure over $U$. Each SCM $\mathcal{M}$ is associated to a causal diagram $\mathcal{G}$ (Bareinboim et al., 2022) over the node set $V$ where $V_{i}\rightarrow V_{j}$ if $V_{i}$ is an argument of $f_{V_{j}}$, and $V_{i}\dashleftarrow\!\!\!\!\!\!\dashrightarrow V_{j}$ if the corresponding $U_{V_{i}},U_{V_{j}}$ are not independent. An instantiation of the exogenous variables $U=u$ is called a unit. By $Y_{x}(u)$ we denote the potential response of $Y$ when setting $X=x$ for the unit $u$, which is the solution for $Y(u)$ to the set of equations obtained by evaluating the unit $u$ in the submodel $\mathcal{M}_{x}$, in which all equations in $\mathcal{F}$ associated with $X$ are replaced by $X=x$. Throughout the paper, we assume a specific cluster causal diagram $\mathcal{G}_{\text{SFM}}$ known as the standard fairness model (SFM) (Plečko and Bareinboim, 2024) over endogenous variables $\{X,Z,W,Y,\widehat{Y}\}$ shown in Fig. 3. The SFM consists of the following: protected attribute, labeled $X$ (e.g., gender, race, religion), assumed to be binary; the set of confounding variables $Z$, which are not causally influenced by the attribute $X$ (e.g., demographic information, zip code); the set of mediator variables $W$ that are possibly causally influenced by the attribute (e.g., educational level or other job-related information); the outcome variable $Y$ (e.g., GPA, salary); the predictor of the outcome $\widehat{Y}$ (e.g., predicted GPA, predicted salary). The SFM encodes the assumptions typically used in the causal inference literature about the lack of hidden confounding. The availability of the SFM and the implied assumptions are a possible limitation of the paper, while we note that partial identification techniques for bounding effects can be used for relaxing them (Zhang et al., 2022).
@@
Figure 2: Standard Fairness Model.
$\emptyset$$\{I\}$$\{D\}$$\{S\}$$\{D,I\}$$\{I,S\}$$\{D,S\}$$\{D,I,S\}$
Figure 3: Graphical representation $\mathcal{G}_{\text{PSEL}}$.
Based on the SFM, we will use the following causal fairness measures:
Definition 1 (Population-level Causal Fairness Measures (Pearl, 2001; Plečko and Bareinboim, 2024)).
@@
A causally-fair predictor for a subset $S$ of the above measures is defined as:
Definition 2 (Causally Fair Predictor (Plecko and Bareinboim, 2024b)).
$\widehat{Y}^{S}=$ $\operatorname*{arg\,min}_{f}$ $\qquad\mathbbm{E}\;$ $L(Y,f(X,Z,W))$ (15)
subject to $\;\;\;\text{NDE}_{x_{0},x_{1}}(f)$ $=\text{NDE}_{x_{0},x_{1}}(y)\cdot\mathbb{1}(\text{DE}\notin S)$ (16)
@@
The definition of $\widehat{Y}^{S}$ has a straightforward interpretation. For any pathway in the set $S$, the corresponding causal effect should be 00, as proposed in the path-specific causal fairness literature (Nabi and Shpitser, 2018; Chiappa, 2019). However, importantly, pathways that are not in $S$ also need to be constrained – the effect of $X$ on $\widehat{Y}$ along these paths should not change compared to the true outcome $Y$ (Plecko and Bareinboim, 2024b). For instance, if the direct path is not in $S$, then we expect to have $\text{NDE}_{x_{0},x_{1}}(\widehat{y})=\text{NDE}_{x_{0},x_{1}}(y)$ (and similarly for IE, SE).
@@ 2 Path-Specific Excess Loss
Let $L(\widehat{Y},Y)$ be a loss function and $\widehat{Y}^{S}$ the optimal $S$-fair predictor with respect to $L$. Define the path-specific excess loss (PSEL) of a pair $S,S^{\prime}$ as:
$\text{PSEL}(S\to S^{\prime})=\mathbbm{E}[L(\widehat{Y}^{S^{\prime}},Y)]-\mathbbm{E}[L(\widehat{Y}^{S},Y)].$ (20)
@@
The total excess loss computes the increase in the loss for the totally constrained predictor $\widehat{Y}^{\{D,I,S\}}$ with direct, indirect, and spurious effects removed compared to the unconstrained predictor $\widehat{Y}^{\emptyset}$.  [^2] Generally, other classifiers that remove only subsets of the causal paths between $X$ and $Y$ may be considered fair, depending on considerations of business necessity (Plečko and Bareinboim, 2024). Still, the rationale developed in this paper can be easily adapted to such settings. The total excess loss can be decomposed as a sum of path-specific excess losses (all proofs are provided in Appendix B):
@@
$\text{PSEL}(\emptyset\to\{D,I,S\})$ $=\text{PSEL}(\emptyset\to\{D\})+\text{PSEL}(\{D\}\to\{D,I\})$ (21)
$\quad+\text{PSEL}(\{D,I\}\to\{D,I,S\}).$ (22)
Remark 2 (Non-Uniquness of Decomposition).
@@
$\text{PSEL}(\emptyset\to\{S_{1}\})+\text{PSEL}(\{S_{1}\}\to\{S_{1},S_{2}\})+\text{PSEL}(\{S_{1},S_{2}\}\to\{D,I,S\})$ (23)
@@
Fig. 3 provides a graphical overview of all the possible path-specific excess loses. In the left side, we start with $S=\emptyset$ and the predictor $\widehat{Y}^{\emptyset}$. Then, we can add any of $\{D,I,S\}$ to the $S$-set, to obtain the predictors $\widehat{Y}^{D},\widehat{Y}^{I}$, or $\widehat{Y}^{S}$, and so on. The graph representing all the possible states $\widehat{Y}^{S}$ and transitions between pairs $(\widehat{Y}^{S},\widehat{Y}^{S^{\prime}})$ shown in Fig. 3 is labeled $\mathcal{G}_{\text{PSEL}}$. There are six paths starting from $\emptyset$ and ending in $\{D,I,S\}$. In Alg. 1, we introduce a procedure that sweeps over all the edges and paths in $\mathcal{G}_{\text{PSEL}}$ to compute path-specific excess losses, while also computing the change in the TV measure between groups in order to track the reduction in discrimination.
@@
6       compute the average path-specific excess loss and TV difference across all paths $\emptyset\to\dots\to\{D,I,S\}$,
$\text{APSEL}(S_{i})$ $=\frac{1}{3!}\sum_{\begin{subarray}{c}\text{paths }\pi\in\mathcal{G}_{\text{PSEL}}:\\ \emptyset\text{ to }\{D,I,S\}\end{subarray}}\text{PSEL}(\pi^{<S_{i}}\to\pi^{<S_{i}}\cup S_{i})$ (24)
$\text{ATVD}(S_{i})$ $=\frac{1}{3!}\sum_{\begin{subarray}{c}\text{paths }\pi\in\mathcal{G}_{\text{PSEL}}:\\ \emptyset\text{ to }\{D,I,S\}\end{subarray}}\text{TVD}(\pi^{<S_{i}}\to\pi^{<S_{i}}\cup S_{i}),$ (25)
@@
$\text{TVD}(S\to S^{\prime})=\underbrace{\mathbbm{E}[Y^{S^{\prime}}\mid x_{1}]-\mathbbm{E}[Y^{S^{\prime}}\mid x_{0}]}_{\text{TV after removing }S^{\prime}\setminus S}-\underbrace{\mathbbm{E}[Y^{S}\mid x_{1}]-\mathbbm{E}[Y^{S}\mid x_{0}]}_{\text{TV before removing }S^{\prime}\setminus S}.$ (26)
The quantities $\text{PSEL}(S\to S^{\prime})$ and $\text{TVD}(S\to S^{\prime})$ are naturally associated with the effect that was removed, i.e., $S^{\prime}\setminus S$. As there are multiple ways of reaching the set $\{D,I,S\}$ from $\emptyset$ in $\mathcal{G}_{\text{PSEL}}$, each of the corresponding effects (direct, indirect, spurious) will be associated with a number of different PSELs and TVDs (generally, note that the complexity is exponential in the number of causal paths included). In Eqs. 24-25 inside the algorithm, we compute the average PSEL and TVD across all the edges that are associated with a specific effect $S_{i}$. This simple intuition, corresponding to taking an average across all of the possible decompositions of the total excess loss (Eq. 23), turns out to be equivalent to a Shapley value (Shapley et al., 1953) of a suitably chosen value function:
@@
$f_{1}(S)$ $=\text{PSEL}(\emptyset\to S),f_{2}(S)=\text{TVD}(\emptyset\to S).$ (27)
The Shapley value for the effect $S_{i}\in\{D,I,S\}$ and function $f_{k}$, is computed as
$\phi^{k}(S_{i})=\sum_{S\subseteq\{D,I,S\}\setminus\{S_{i}\}}\frac{1}{n\binom{n-1}{|S|}}\left(f_{k}(S\cup\{S_{i}\})-f_{k}(S)\right).$ (28)
@@
$\phi^{1}(S_{i})$ $=\text{APSEL}(S_{i}),\phi^{2}(S_{i})=\text{ATVD}(S_{i}),$ (29)
with $\text{APSEL}(S_{i}),\text{ATVD}(S_{i})$ defined in Eqs. 24, 25, respectively.
The above proposition illustrates how averaging the influence of removing a causal effect over all possible ways of reaching $\{D,I,S\}$ from $\emptyset$ is equivalent to computing the Shapley values of an appropriate value function $f$. We next introduce the notion of a causal fairness/utility ratios:
@@ 3 Causally-Fair Constrained Learning
In the preceding section, we developed an approach for quantifying the tension between fairness and accuracy from a causal viewpoint. The results were contingent on finding the optimal causally-fair predictors $\widehat{Y}^{S}$ following Def. 2. However, computing the predictors $\widehat{Y}^{S}$ is quite challenging in practice, due to several complex causal constraints in the optimization problem. In this sectionm we develop a practical approach for solving this problem. We begin by introducing a Lagrangian form of the optimal causally-fair predictor:
Input: training data $\mathcal{D}_{t}$, evaluation data $\mathcal{D}_{e}$, precision $\epsilon$

### v2

@@ 1.2 Preliminaries
Figure 2: Standard Fairness Model, with the protected attribute $X$, set of confounders $Z$, set of mediators $W$, outcome $Y$, and a predictor $\widehat{Y}$.
We use the language of structural causal models (SCMs) (Pearl 2000). An SCM is a tuple $\mathcal{M}:=\langle V,U,\mathcal{F},P(u)\rangle$ , where $V$, $U$ are sets of endogenous (observable) and exogenous (latent) variables, respectively, $\mathcal{F}$ is a set of functions $f_{V_{i}}$, one for each $V_{i}\in V$, where $V_{i}\leftarrow f_{V_{i}}(\mathrm{pa}(V_{i}),U_{V_{i}})$ for some $\mathrm{pa}(V_{i})\subseteq V$ and $U_{V_{i}}\subseteq U$. $P(u)$ is a strictly positive probability measure over $U$. Each SCM $\mathcal{M}$ is associated to a causal diagram $\mathcal{G}$ (Bareinboim et al. 2022) over the node set $V$ where $V_{i}\rightarrow V_{j}$ if $V_{i}$ is an argument of $f_{V_{j}}$, and $V_{i}\dashleftarrow\!\!\!\!\!\!\dashrightarrow V_{j}$ if the corresponding $U_{V_{i}},U_{V_{j}}$ are not independent. An instantiation of the exogenous variables $U=u$ is called a unit. By $Y_{x}(u)$ we denote the potential response of $Y$ when setting $X=x$ for the unit $u$, which is the solution for $Y(u)$ to the set of equations obtained by evaluating the unit $u$ in the submodel $\mathcal{M}_{x}$, in which all equations in $\mathcal{F}$ associated with $X$ are replaced by $X=x$. For more details about the causal inference background, we refer the reader to (Pearl 2000; Bareinboim et al. 2022; Plečko and Bareinboim 2024). Throughout the paper, we assume a specific cluster causal diagram $\mathcal{G}_{\text{SFM}}$ known as the standard fairness model (SFM) (Plečko and Bareinboim 2024) over endogenous variables $\{X,Z,W,Y,\widehat{Y}\}$ shown in Fig. 2. The SFM consists of the following: protected attribute, labeled $X$ (e.g., gender, race, religion), assumed to be binary; the set of confounding variables $Z$, which are not causally influenced by the attribute $X$ (e.g., demographic information, zip code); the set of mediator variables $W$ that are possibly causally influenced by the attribute (e.g., educational level or other job-related information); the outcome variable $Y$ (e.g., GPA, salary); the predictor of the outcome $\widehat{Y}$ (e.g., predicted GPA, predicted salary). The SFM encodes the assumptions typically used in the causal inference literature about the lack of hidden confounding. The availability of the SFM and the implied assumptions are a possible limitation of the paper, while we note that partial identification techniques for bounding effects can be used for relaxing them (Zhang, Tian, and Bareinboim 2022). Based on the SFM, we will use the following causal fairness measures:
Definition 1 (Population-level Causal Fairness Measures (Pearl 2001; Plečko and Bareinboim 2024)).
@@
The NDE in Eq. 26 compares the potential outcome $Y_{x_{1},W_{x_{0}}}$, in which $Y$ responds to $X=x_{1}$ along the direct path, while $W$ is set at the value it would attain naturally when responding to $X=x_{0}$, against the potential outcome $Y_{x_{0}}$ where $X=x_{0}$ along both direct and indirect paths. In this way, the NDE measures the variations induced by changing $x_{0}\to x_{1}$ along the direct causal path, quantifying direct discrimination. Similarly, the NIE in Eq. 27 compares $Y_{x_{1},W_{x_{0}}}$ vs. $Y_{x_{1}}$, and thus captures variations induced by considering a change $x_{1}\to x_{0}$ along the indirect causal path (note that both $Y_{x_{1},W_{x_{0}}}$ and $Y_{x_{1}}$ respond to $X=x_{1}$ along the direct path, so only indirect variations are induced when taking the difference). Finally, the NSE in Eq. 28 compares $Y\mid X=x$ vs. $Y_{x}$. In the former, due to conditioning on $X=x$, the distribution over the set of confounders $Z$ (in Fig. 2) changes according to this conditioning, while in the potential outcome $Y_{x}$, the distribution of $Z$ does not change, since $X=x$ is set by intervention. Therefore, taking the difference captures the spurious effect of $X$ on $Y$, along the backdoor path $X\dashleftarrow\!\!\!\!\!\!\dashrightarrow Z\to Y$. A causally-fair predictor for a subset $S$ of the above measures is defined as:
Definition 2 (Causally Fair Predictor (Plecko and Bareinboim 2024b)).
The optimal causally $S$-fair predictor $\widehat{Y}^{S}$ with respect to a loss function $L$ and pathways in $S$ is the solution to the following optimization problem:
$\operatorname*{arg\,min}_{f}$ $\quad\mathds{E}\;$ $L(Y,f(X,Z,W))$ (29)
s.t. $\;\text{NDE}_{x_{0},x_{1}}(f)$ $=\text{NDE}_{x_{0},x_{1}}(y)\cdot\mathbb{1}(\text{DE}\notin S)$ (30)
@@
The definition of $\widehat{Y}^{S}$ has a straightforward interpretation. For any pathway in the set $S$, the corresponding causal effect should be 00, as proposed in the path-specific causal fairness literature (Nabi and Shpitser 2018; Chiappa 2019). However, importantly, pathways that are not in $S$ also need to be constrained – the effect of $X$ on $\widehat{Y}$ along these paths should not change compared to the true outcome $Y$ (Plecko and Bareinboim 2024b). For instance, if the direct path is not in $S$ (meaning it is considered to be non-discriminatory), then we expect to have $\text{NDE}_{x_{0},x_{1}}(\widehat{y})=\text{NDE}_{x_{0},x_{1}}(y)$ (and similarly for IE, SE). This form of constraint on discriminatory pathways ensures that no undesirable bias amplification occurs along non-discriminatory pathways.
@@ 2 Path-Specific Excess Loss
Let $L(\widehat{Y},Y)$ be a loss function and $\widehat{Y}^{S}$ the optimal causally $S$-fair predictor with respect to $L$. Define the path-specific excess loss (PSEL) of a pair $S,S^{\prime}$ as:
$\text{PSEL}(S\to S^{\prime})=\mathds{E}[L(\widehat{Y}^{S^{\prime}},Y)]-\mathds{E}[L(\widehat{Y}^{S},Y)].$ (34)
@@
The total excess loss computes the increase in the loss for the totally constrained predictor $\widehat{Y}^{\{D,I,S\}}$ with direct, indirect, and spurious effects removed compared to the unconstrained predictor $\widehat{Y}^{\emptyset}$.  [^1] Other classifiers that remove only subsets of the causal paths between $X$ and $Y$ may be considered fair, depending on considerations of business necessity (Plečko and Bareinboim 2024). The rationale developed in this paper can be adapted to such settings. In the following theorem, we show that the total excess loss can be decomposed as a sum of path-specific excess losse. All proofs are provided in Appendix B (for supplements, see full paper version at https://arxiv.org/abs/2405.15443):
@@
$\text{PSEL}(\emptyset\to\{D,I,S\})$ $=\text{PSEL}(\emptyset\to\{D\})$ (35)
$+\text{PSEL}(\{D\}\to\{D,I\})$ (36)
$+\text{PSEL}(\{D,I\}\to\{D,I,S\}).$ (37)
Remark 2 (Non-Uniqueness of Decomposition).
@@
$\text{PSEL}(\emptyset\to\{S_{1}\})+\text{PSEL}(\{S_{1}\}\to\{S_{1},S_{2}\})$ (38)
$\quad+\text{PSEL}(\{S_{1},S_{2}\}\to\{D,I,S\})$ (39)
@@
$\emptyset$$\{I\}$$\{D\}$$\{S\}$$\{D,I\}$$\{I,S\}$$\{D,S\}$$\{D,I,S\}$
Figure 3: Graphical representation $\mathcal{G}_{\text{PSEL}}$.
Fig. 3 provides a graphical overview of all the possible path-specific excess loses. In the left side, we start with $S=\emptyset$ and the predictor $\widehat{Y}^{\emptyset}$. Then, we can add any of $\{D,I,S\}$ to the $S$-set, to obtain the predictors $\widehat{Y}^{D},\widehat{Y}^{I}$, or $\widehat{Y}^{S}$, and so on. The graph representing all the possible states $\widehat{Y}^{S}$ and transitions between pairs $(\widehat{Y}^{S},\widehat{Y}^{S^{\prime}})$ shown in Fig. 3 is labeled $\mathcal{G}_{\text{PSEL}}$. There are six paths starting from $\emptyset$ and ending in $\{D,I,S\}$. In Alg. 1, we introduce a procedure that sweeps over all the edges and paths in $\mathcal{G}_{\text{PSEL}}$ to compute path-specific excess losses, while also computing the change in the TV measure between groups in order to track the reduction in discrimination. In the main body of the paper we discuss fairness-utility trade-offs when considering direct, indirect, and spurious effects. In Appendix F we described how to adapt all the results to a general setting considering more granular path-specific effects.
@@
6       compute the average path-specific excess loss and TV difference across all paths $\emptyset\to\dots\to\{D,I,S\}$ in $\mathcal{G}_{\text{PSEL}}$
7       $\text{APSEL}(S_{i})$ is computed as
$\frac{1}{3!}\sum_{\begin{subarray}{c}\pi\in\mathcal{G}_{\text{PSEL}}:\\ \emptyset\text{ to }\{D,I,S\}\end{subarray}}\text{PSEL}(\pi^{<S_{i}}\to\pi^{<S_{i}}\cup S_{i})$ (40)
8       $\text{ATVD}(S_{i})$ is computed as
$\frac{1}{3!}\sum_{\begin{subarray}{c}\pi\in\mathcal{G}_{\text{PSEL}}:\\ \emptyset\text{ to }\{D,I,S\}\end{subarray}}\text{TVD}(\pi^{<S_{i}}\to\pi^{<S_{i}}\cup S_{i}),$ (41)
@@ 8       $\text{ATVD}(S_{i})$ is computed as
$\text{TVD}(S\to S^{\prime})=$ $\underbrace{\mathds{E}[Y^{S^{\prime}}\mid x_{1}]-\mathds{E}[Y^{S^{\prime}}\mid x_{0}]}_{\text{TV after removing }S^{\prime}\setminus S}$ (42)
$-\underbrace{\mathds{E}[Y^{S}\mid x_{1}]-\mathds{E}[Y^{S}\mid x_{0}]}_{\text{TV before removing }S^{\prime}\setminus S}.$ (43)
The quantities $\text{PSEL}(S\to S^{\prime})$ and $\text{TVD}(S\to S^{\prime})$ are naturally associated with the effect that was removed, i.e., $S^{\prime}\setminus S$. In this context, we mention a connection with previous work (Zhang and Bareinboim 2018b), which provides a way of quantifying direct, indirect, and spurious effects of the attribute $X$ on the outcome $Y$. In Appendix C, we show that our approach of quantifying the change in the TV measure through the TVD quantity in practice closely corresponds the approach proposed by (Zhang and Bareinboim 2018b).
As there are multiple ways of reaching the set $\{D,I,S\}$ from $\emptyset$ in $\mathcal{G}_{\text{PSEL}}$, each of the corresponding effects (direct, indirect, spurious) will be associated with a number of different PSELs and TVDs (generally, note that the complexity is exponential in the number of causal paths included). In Eqs. 40-41 inside the algorithm, we compute the average PSEL and TVD across all the edges that are associated with a specific effect $S_{i}$. This simple intuition, corresponding to taking an average across all of the possible decompositions of the total excess loss (Eq. 38), turns out to be equivalent to a Shapley value (Shapley et al. 1953) of a suitably chosen value function:
@@
$f_{1}(S)$ $=\text{PSEL}(\emptyset\to S)$ (44)
$f_{2}(S)$ $=\text{TVD}(\emptyset\to S).$ (45)
The Shapley value $\phi^{k}(S_{i})$ for the effect $S_{i}\in\{D,I,S\}$ and function $f_{k}$, is computed as
$\sum_{S\subseteq\{D,I,S\}\setminus\{S_{i}\}}\frac{1}{n\binom{n-1}{|S|}}\left(f_{k}(S\cup\{S_{i}\})-f_{k}(S)\right).$ (46)
@@
$\phi^{1}(S_{i})$ $=\text{APSEL}(S_{i})$ (47)
$\phi^{2}(S_{i})$ $=\text{ATVD}(S_{i}),$ (48)
with $\text{APSEL}(S_{i})$ defined as
$\frac{1}{3!}\sum_{\begin{subarray}{c}\pi\in\mathcal{G}_{\text{PSEL}}:\\ \emptyset\text{ to }\{D,I,S\}\end{subarray}}\text{PSEL}(\pi^{<S_{i}}\to\pi^{<S_{i}}\cup S_{i}),$ (49)
and $\text{ATVD}(S_{i})$ defined as
$\frac{1}{3!}\sum_{\begin{subarray}{c}\pi\in\mathcal{G}_{\text{PSEL}}:\\ \emptyset\text{ to }\{D,I,S\}\end{subarray}}\text{TVD}(\pi^{<S_{i}}\to\pi^{<S_{i}}\cup S_{i}).$ (50)
The above proposition illustrates how averaging the influence of removing a causal effect over all possible ways of reaching $\{D,I,S\}$ from $\emptyset$ is equivalent to computing the Shapley values of an appropriate value function $f$. We remark that the attribution in Prop. 3 is not the only approach one could take. An alternative would be to average the contributions of each pathways across all edges of the graph $\mathcal{G}_{\text{PSEL}}$ (instead of focusing on all pathways from $\emptyset$ to $\{D,I,S\}$). Such an attribution would not satisfy the Shapley axioms, however. We next introduce the notion of a causal fairness/utility ratios:
@@
Whenever $\text{APSEL}(S_{i})=0$, $\text{CFUR}(S_{i})$ is equal to 00.
@@ 3 Causally-Fair Constrained Learning
In the preceding section, we developed an approach for quantifying the tension between fairness and accuracy from a causal viewpoint. The results were contingent on finding the optimal causally-fair predictors $\widehat{Y}^{S}$ following Def. 2. However, computing the predictors $\widehat{Y}^{S}$ is quite challenging in practice, due to several complex causal constraints in the optimization problem. We now develop a practical approach for solving this problem, by first introducing a Lagrangian form of the optimal causally-fair predictor:
Definition 5 (Causal Lagrange Predictor $\widehat{Y}^{S}$).
The causally $S$-fair $\lambda$-optimal predictor $\widehat{Y}^{S}(\lambda)$ with respect to pathways in $S$ and the loss function $L$ is the solution of the following:
$\operatorname*{arg\,min}_{f}$ $\;\mathds{E}\;L(Y,f(X,Z,W))+$ (52)
$\lambda\big{(}\text{NDE}_{x_{0},x_{1}}(f)-\text{NDE}_{x_{0},x_{1}}(y)\cdot\mathbb{1}(\text{DE}\notin S)\big{)}^{2}+$ (53)
$\lambda\big{(}\text{NSE}_{x_{0}}(f)-\text{NSE}_{x_{0}}(y)\cdot\mathbb{1}(\text{SE}\notin S)\big{)}^{2}+$ (55)
$\lambda\big{(}\text{NSE}_{x_{1}}(f)-\text{NSE}_{x_{1}}(y)\cdot\mathbb{1}(\text{SE}\notin S)\big{)}^{2}$ (56)
Input: training data $\mathcal{D}_{t}$, evaluation data $\mathcal{D}_{e}$, set $S$, precision $\epsilon$